OpenStudy (anonymous):

can someone please give me h,k and the radius for this equation of a circle below.

OpenStudy (anonymous):

\[5x ^{2} -9x +5y ^{2} -9x -2 =0\]

OpenStudy (imstuck):

You have to complete the square to do that. Do you know how?

OpenStudy (imstuck):

I think you may have typed that wrong. You have two "-9x"'s and no y term.

OpenStudy (mathstudent55):

You wrote: \(5x ^{2} -9x +5y ^{2} -9\color{red}{x} -2 =0\) Is it supposed to be: \(5x ^{2} -9x +5y ^{2} -9\color{red}{y} -2 =0\)

OpenStudy (anonymous):

i know that we need to pull the 5 out and complete the square problem i am having is figuring out how o complete the square. i have done many problems like this. i am just very confused and frustrated and i am trying to figure out why i am messing up on this problem. can you please show me what to do regarding this problem.

OpenStudy (anonymous):

and yes you are right @mathstudent55, my bad i was typing fast

OpenStudy (imstuck):

Sure I can show you!

OpenStudy (imstuck):